OpenStudy (anonymous):

Which statement best explains the value of 16 − (−3)? The additive inverse of −3 is −3, so 16 − (−3) = 19. The additive inverse of −3 is −3, so 16 − (−3) = 13. The additive inverse of −3 is +3, so 16 − (−3) = 19. The additive inverse of −3 is +3, so 16 − (−3) = 13.

OpenStudy (anonymous):

who knows pretty clear that \(16-(-3)=16+3=9\) though

OpenStudy (anonymous):

I have no idea @NKP

OpenStudy (anonymous):

go with C it is always C

OpenStudy (anonymous):

are you guessing @satellite73

OpenStudy (anonymous):

yes but it is an educated guess

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since both statements are correct